OpenStudy (anonymous):

Derivative...Chain Rule (4 + cosx)^6 I see chain rule, do you ?

OpenStudy (anonymous):

I got 6(4 + cosx)(-sinx) but I'm missing something! @SolomonZelman

OpenStudy (anonymous):

@mathaddict4471

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (solomonzelman):

\[\frac{d}{dx}(4 + \cos x)^6=6(4 + \cos x)^{6-1}\times (-\sin x),\]

OpenStudy (anonymous):

whoops!! the 6-1 ah thank you!

OpenStudy (anonymous):

so it would be -6(sinx)(4 +cosx)^5

OpenStudy (solomonzelman):

yes

jimthompson5910 (jim_thompson5910):

Outer function: f(x) = x^6 f ' (x) = 6x^5 Inner function: g(x) = 4+cos(x) g ' (x) = -sin(x) Composite function: h(x) = f(g(x)) = [4 + cos(x)]^6 chain rule: h ' (x) = f ' (g(x)) * g ' (x) but it looks like you already got it

OpenStudy (anonymous):

thank you @jim_thompson5910 =) every bit helps !

OpenStudy (anonymous):

@SolomonZelman I am terrible at u substition Still struggling, what gives ?

OpenStudy (solomonzelman):

You just have to understand this: (when f(x) is the inner function) u = f(x) du = f'(x) dx

OpenStudy (anonymous):

okay I am grasping that part the best its just after I find the u and du/dx I am lost

OpenStudy (anonymous):

i am reading this right now http://www.mathcentre.ac.uk/resources/workbooks/mathcentre/web-integrationbysub-tony.pdf